Count the atoms of each element in

CH3​(CH2​)9​O2

​



1
Expert's answer
2021-04-05T07:42:29-0400

Q177318


Count the atoms of each element in

CH3​(CH2​)9​O2.


Solution:


Let us expand the given structure and then find the number of atoms of each element.



"CH_3-CH_2-CH_2-CH_2-CH_2-CH_2-CH_2-CH_2-CH_2-CH_2-O_2"


count the # of carbon atoms first.

# of carbon = 10.


Next count the # of Hydrogen

# of hydrogen atom

= 3 + 2 + 2+ 2 + 2 + 2 + 2 + 2 + 2 + 2

= 3 + 9 ( 2 )

= 3 + 18

= 21 hydrogen atoms.


Now count the # of oxygen atom.

There is a subscript 2 present on O.

which means there is 2 atoms of O.

So, # of oxygen atom = 2.



# of carbon atoms = 10

# of hydrogen atoms = 21.

# of oxygen atoms = 2.